OpenStudy (anonymous):

can somebody explain how to find the formula to the inverse function log2 6-log2 15+log 20

OpenStudy (anonymous):

do you mean the inverse variation?

OpenStudy (anonymous):

sorry find the exact value of the expression

OpenStudy (anonymous):

do you agree that log(x^(2)) = 2log(x) ? and that x^(-1) = 1/x

OpenStudy (anonymous):

yes the one to one value ?

OpenStudy (anonymous):

log_2(6)-log_2(15)+log(20) so log_2(6) + log_2(15^(-1)) + log(20) = log_2(6/15) + log20

OpenStudy (anonymous):

you can't simplify it anymore because the the logrithims don't share the same base

OpenStudy (anonymous):

*logarithms

OpenStudy (anonymous):

do you want to find the inverse of this? anyfeed back would be appreciated. BTW you can't take log(0) because A^(x) =/= 0 A being any real number

OpenStudy (anonymous):

what ? i thought I was suppose to divide do log2 6*20 and divide by 15...

OpenStudy (campbell_st):

thew question is \[\log _{2} (6/15 \times20) = \log _{2}(8) = 3\] this is becuase 2^3 = 8

OpenStudy (anonymous):

aah yes

OpenStudy (anonymous):

campbell_st wth?

OpenStudy (anonymous):

repost the question so it isn't so werid

OpenStudy (anonymous):

you can't multiply by 20 as the log is not of the same base thus they cannot be combined cambell

OpenStudy (campbell_st):

well I assumed that it was base 2.... if base 10 then use change of base on all...to say base e

OpenStudy (anonymous):

you shouldn't assume things when it comes to this kind of math but yeah if he wrote the question with all log_2 log_2(x) = y x = 2^(y) then it would be log_2((6/15)*20) log_2(8) = y so 8 = 2^(y) y = 3

OpenStudy (anonymous):

therefore, log2 6-log2 15+log2 20 = 3